Suppose cosθ= 2/5 and θ is in quadrant IV.
Use the Fundamental Trigonometric Identities to find sin θ.

Answer by stanbon(75887)

Suppose cosθ= 2/5 and θ is in quadrant IV.
Use the Fundamental Trigonometric Identities to find sin θ.
------
Note: x is positive and y is negative in QIV.
By definition cos(t) = x/r
So, x = 2 and r = 5
----
Then y = -sqrt[5^2-2^2] = -sqrt(21)
----
Ans: sin(t) = y/r = -sqrt(21)/5
-------------
Cheers,
Stan H.
----------